The sum of two numbers is 90. The larger number is 14 more than 3 times the smaller number. Find the numbers.
Answer by Hawksfan(61)

let x = larger number
y = smaller


x = 3y + 14
x + y = 90
(3y + 14) + y = 90
4y + 14 = 90
subtract 14
4y = 76
y = 19
x = 71